7. Consider three five-year Treasury bonds at par ($1,000) with coupon of 2%. 4%, and 6% respectively.

(a) Compute their durations. (b) What can you conclude about the relationship between duration and coupon?
Answer #1

Duration of 2% bond=(1*2%*1000/1.02+2*2%*1000/1.02^2+3*(1000+2%*1000)/1.02^3)/1000=2.941560938

Duration of 4% bond=(1*4%*1000/1.04+2*4%*1000/1.04^2+3*(1000+4%*1000)/1.04^3)/1000=2.886094675

Duration of 6% bond=(1*6%*1000/1.06+2*6%*1000/1.06^2+3*(1000+6%*1000)/1.06^3)/1000=2.833392666

Duration is inversely proportional to coupon